News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 19 January 2007 8.30am EST

During the last week Chandra completed the observing schedule as planned.

A real-time procedure was run on 16 Jan to raise the Aspect Camera Assembly (ACA) CCD temperature from -20 deg C back to -19 deg C. This step was taken after monitoring of the CCD temperature following the decrease to -20 deg C on 26 Dec indicated that the new temperature provided insufficient control margin of the CCD Thermal Electric Cooler (TEC). The performance of the CCD is expected to be acceptable at -19 deg C. Note that contrary to last week's report, the ACA dark current calibration conducted on a 5 Jan indicated the expected decrease (not increase) in warm pixels due to the lowering of the CCD temperature.

A Chandra image release was issued on 9 Jan of the Kepler Supernova Remnant. The spectacular image was acquired over 750Ks and the data indicate that Kepler could be the nearest example of a relatively rare "prompt" Type Ia explosion, which occur in more massive progenitors only about 100 million years after the star formed rather than several billion years. For details see: http://chandra.harvard.edu/photo/2007/kepler/

The schedule of targets for next week is shown below.

------------------------------------------
Radiation Belts                      Jan 21
GRB 061121              ACIS-S       Jan 22
NGC 1893                ACIS-I
Cyg X-2                 HRC-I        Jan 24
HCG 100                 ACIS-I
Radiation Belts
PKS0745-191             ACIS-I
PKS 1055+201            ACIS-S       Jan 25
Radiation Belts                      Jan 27
PKS 1055+201            ACIS-S
HE1113-0641             ACIS-S       Jan 28
1RXS J171824.2-40293    HRC-S
ZwCl1742                HRC-I

------------------------------------------

All spacecraft subsystems continued to support nominal operations.

Return to Status Reports